A crazy physics teacher has a golden spear of 3.2m. How long it will be when he heats it up to 1000C from room temperature(+20C)?

Single choice.






1
Expert's answer
2021-06-01T16:24:21-0400
"L'=L_0(1+\alpha\Delta T)\\L'=3.2(1+14\cdot10^{-6}(1000-20))=3.244\ m"